### v1

# A Visual-Inertial Motion Prior SLAM for Dynamic Environments

Abstract
The Visual-Inertial Simultaneous Localization and Mapping (VI-SLAM) algorithms which are mostly based on static assumption are widely used in fields such as robotics, UAVs, VR, and autonomous driving. To overcome the localization risks caused by dynamic landmarks in most VI-SLAM systems, a robust visual-inertial motion prior SLAM system, named (IDY-VINS), is proposed in this paper which effectively handles dynamic landmarks using inertial motion prior for dynamic environments to varying degrees. Specifically, potential dynamic landmarks are preprocessed during the feature tracking phase by the probabilistic model of landmarks’ minimum projection errors which are obtained from inertial motion prior and epipolar constraint. Subsequently, a bundle adjustment (BA) residual is proposed considering the minimum projection error prior for dynamic candidate landmarks. This residual is integrated into a sliding window based nonlinear optimization process to estimate camera poses, IMU states and landmark positions while minimizing the impact of dynamic candidate landmarks that deviate from the motion prior. Finally, experimental results demonstrate that our proposed system outperforms state-of-the-art methods in terms of localization accuracy and time cost by robustly mitigating the influence of dynamic landmarks.
I INTRODUCTION
In environments without satellite positioning, V-SLAM obtains information of the surrounding environment through visual sensors to solve its own motion. It has the advantages of low cost, high precision in trajectory estimation within small areas, no need for any environmental prior, and rich applicable scenarios. However, visual-only sensor is easily affected by external environmental and motion factors, such as dynamic landmarks, the lack of texture, lighting conditions, motion blur, and few common field of view (FOV) [1]. Therefore, an inertial measurement unit (IMU), which can provide relatively accurate pose estimation information without being affected by the environment in a short period of time, can be integrated to compensate for the deficiencies of visual-only sensor [2].
However, the assumption of static landmarks in visual-related SLAM still makes it has hidden localization risk in environments with dynamic landmarks [3]. The dynamic landmarks will affect the accuracy of visual positioning, and in extreme cases, they will lead to rapid deterioration of positioning.
In recent years, research on visual-related SLAM in dynamic environments has gradually become a focus. Most studies seek solutions through deep learning methods such as semantic recognition [4, 5] while some studies use geometric methods [6, 7]. However, deep learning methods are always limited by scenes and training processes while geometric methods always have low-precision and unstable performance.
In order to solve the problem of dynamic landmarks, a robust visual-inertial motion prior SLAM system is proposed in this paper which is applicable to a large range of dynamic environments and only uses the motion prior propagated by the IMU in a short period of time to process dynamic landmarks. This paper presents our work from the following aspects:
(1) The preprocessing of potential dynamic landmarks with inertial motion prior: The minimum projection errors of landmarks are obtained from inertial motion prior and epipolar constraint to eliminate dynamic landmarks that significantly deviate from the motion prior and mark dynamic candidate landmarks during the feature tracking phase.
(2) The bundle adjustment (BA) residual for dynamic candidate landmarks: A dynamic visual residual is constructed which is used in nonlinear optimization process based on sliding window to reduce the impact of landmarks that deviate from the motion prior.
Compared with existing methods in the VIODE dataset [8] which contains dynamic objects within different ranges, superior robust and stable localization accuracy, and good real-time performance of our approach are demonstrated. In addition, the localization accuracy is compared in the EUROC dataset [9], and the result demonstrates stable and superior performance of our approach in static environments.
II RELATED WORK
VI-SLAM algorithms utilize the inertial measurement unit to correct the scale and the camera poses while V-SLAM algorithms only use visual sensors. Generally, they can be divided into filter-based methods and optimization-based methods which have become mainstream. For example, the ROVIO [10] employed the extended Kalman filter (EKF) method, while OKVIS [11] proposed a keyframe-based framework using optimization process. The ORB descriptor [12] for feature matching was employed to optimize camera poses and landmark positions with IMU measurements in ORB-SLAM3 [13]. The VINS-Fusion [14] is an extended version of VINS-Mono [15], which supports stereo cameras with IMU and adopts feature tracking instead of descriptor matching. However, the above VI-SLAM methods still have potential limitations in dealing with dominant dynamic objects.
Recently, many researchers have proposed a variety of methods for dealing with dynamic objects in the V-SLAM and VI-SLAM algorithms.
II-A Methods based on vision geometry
Fan et al. [16] proposed a multi-view geometry-based method using RGB-D camera. The type of each landmark is determined to be dynamic or static according to the geometric relationship between the camera motion and the landmark’s states after obtaining the camera pose by minimizing the reprojection error. The RANSAC method [17, 18] can also be used to remove dynamic landmarks in small-scale dynamic scenes. In fact, it can distinguish between inliers and outliers in various anomalies such as noise, occlusion and dynamic objects, improving the robustness of model parameter estimation. The multi-view geometry-based methods assume that the camera pose estimation is accurate enough, which will lead to the failure of the algorithms when the estimation is inaccurate due to the presence of too many dynamic objects.
II-B Methods based on deep learning
In DynaSLAM [19], the dynamic landmarks in masked areas where predefined dynamic objects by deep learning networks are removed, while the types of remaining landmarks are determined through multi-view geometry. Wen et al. [20] proposed a method that combines semantic segmentation information and spatial motion information of associated pixels to deal with dynamic objects. Masoud S. Bahraini et al. [21] adopted a method that combines multi-level RANSAC operators with deep learning to distinguish between dynamic landmarks and static landmarks. Although deep learning methods can successfully eliminate the landmarks of dynamic objects, there are still some problems, such as: many methods require predefined dynamic objects; when only a part of a dynamic object can be seen due to occlusion, it may not be detected; the detected dynamic objects may not be in movement but are temporarily stationary.
II-C Methods based on inertial sensor
Cui et al. [22] detected and removed dynamic landmarks according to the epipolar constraint provided by the prior of IMU pre-integration, and conducted states estimation by utilizing more accurate static landmarks. The DynaVINS [23] proposed a robust bundle adjustment system based on the IMU pre-integration regularization factor and the weight momentum factor to reject dynamic landmarks. The method in this paper is also based on the motion prior propagated by the IMU which is used to label potential dynamic landmarks, and we propose a robust bundle adjustment residual for dynamic candidate landmarks.
III PREPROCESSING
In this section, the method for preprocessing potential dynamic landmarks with inertial motion prior is introduced. Due to the IMU propagation motion prior between two frames of camera, the minimum projection error is obtained through the landmark’s precision of the epipolar constraint, then its probability model is modeled and applied to determine dynamic landmarks will be eliminated and dynamic candidate landmarks will be marked during the tracking phase.
III-A Inertial Motion Piror
The relative translation and rotation in the camera coordinate system can be calculated by IMU propagation result and the extrinsic parameters between IMU and camera. The residual covariance matrix of IMU pre-integration [24] can be linearly propagated with the first-order approximation, it will be recursively obtained from the initial covariance and the diagonal covariance of noise .
, represents the extrinsic parameters, and respectively are the translation vector and rotation matrix obtained from the propagation in the IMU coordinate system, denotes the IMU coordinate system while denotes the camera coordinate system. represents the IMU between two frames of camera, and is the partial derivative of the residual term and the noise term when the residual is expanded by the first-order Taylor series.
III-B Minimum Projection Error and Preprocessing
Considering the epipolar constraint and the projection relationship of the landmarks between two frames of camera as Fig.2, and under the influence of the two-dimensions’ projection residual, the precision of the epipolar constraint is not exactly zero.
represents the precision of the epipolar constraint, is the coordinate in the normalized camera coordinate system while is the coordinate in the pixel coordinate system, is the coordinate measurement combined the true value and the residual as Fig.2 (b), is the internal parameter matrix of camera.
The minimum value of the square projection error is obtained based on the precision of the epipolar constraint and the residual covariance of IMU pre-integration.
is the minimum square projection error defined by us. is a factor that characterizes the credibility of the motion prior obtained through IMU propagation, which is defined to avoid the incorrect elimination of dynamic landmarks caused by large IMU propagation error (mainly considered as translation error) or the degradation of the epipolar constraint due to an overly small translation amount. is the variance of the position error of IMU pre-integration. Noted that for the pinhole camera model, we adopt the projection error on the pixel plane, while for the fisheye camera model, we use the projection unit length error on the normalized plane.
The square root of projection error of landmark tracked between two frames of camera is modeled with a Gaussian distribution. So in can be interpreted from the probabilistic perspective as the variance of the Gaussian distribution, . Then the distribution of will be approximate to a chi-square distribution with one degree of freedom. The expectation of its coefficient of variation (CV) should approach to . The elimination and marking strategies for potential dynamic landmarks are as follows.
(1) Elimination in preprocess: In the situation where both dynamic and static landmarks exist, the value of CV is always greater than . Sort all landmarks in descending order according to the value of , and start eliminating from the maximum value until the CV no longer decreases or the CV decreases to the target value . The elimination criterion and the recursive formulas for mean, variance and CV of remaining landmarks are as follows.
is the set of landmarks that satisfy above standard. Noted that we implement a strict elimination criterion in order to prevent excessive elimination, only the landmarks with a minimum projection error greater than will execute the elimination operation. is the mean and variance of the remaining landmarks.
(2) Marking in preprocess: Ignoring the influence of large IMU propagation error and the degradation of the epipolar constraint, the remaining landmarks with a minimum projection error () greater than are marked as dynamic candidate landmarks. During the nonlinear optimization process, these landmarks will be treated differently from static landmarks.
(3) Mask for extracting landmarks: After elimination and marking, the area within around the eliminated dynamic landmark is considered the dynamic candidate area and no new landmarks will be extracted.
IV ROBUST BUNDLE ADJUSTMENT
After removing dynamic landmarks and marking dynamic candidate landmarks, tightly coupled [25], high-precision and robust estimation for the camera poses, landmarks and IMU states in the nonlinear optimization process based on sliding window [15] is executed with the goal of minimizing a global residual. And the visual residual for dynamic candidate landmarks is proposed in this paper to further reduce the influence of them on the optimization results considering the role of kernel function.
When a dynamic candidate landmark enters the sliding window based optimization process, it is continuously observed for times. The visual residual for these observations is as which considers the prior of minimum projection error which will also be updated during the iterative optimization process. From the probabilistic perspective, the two-dimensions’ projection residual of dynamic candidate landmarks satisfies a Gaussian distribution with a covariance matrix of minimum value as .
represents the collection of dynamic candidate landmarks in the sliding window, represents the observation to landmark , represents the camera pose state, IMU state and landmarks state that are to be optimized. and is the reduction factor of dynamic candidate landmarks’ visual residuals, and it is related to the prior of minimum projection error (more than one ) obtained through the epipolar constraint precision between the first observation and the observation in the sliding window. is huber kernel function [26]. Noted that .
Combined with the visual residuals of static landmarks, the measurement residuals of the IMU and the prior residual, the global residual to be minimized is (8).
is the collection of static landmarks in the sliding window, and is the covariance matrix of static landmark’s projection residual. represents the collection of inertial measurements while is the residual covariance matrix of the IMU pre-integration [24] between the and frames. represents prior information.
After optimization, 3D-2D projection from the first observation to other relevant observations is performed to calculate the reprojection errors of dynamic candidate landmarks. Remove them outside the sliding window and mark them as confirmed dynamic landmarks if the errors are still relatively large.
V EXPERIMENTAL RESULTS
To evaluate our proposed algorithm, comparisons with the SOTA algorithms in the VIODE and EUROC datasets are carried out, such as VINS-FUSION [14], VINS-FUSION_ransac and Dyna-VINS [23] in the stereo mode.
V-A Datasets
VIODE: VIODE dataset [8] is a simulated dataset that contains lots of dynamic objects, such as cars and trucks. In addition, the dataset includes general occlusion situations. Note that the sub-sequence name none to high means the number and visual field of dynamic objects in the scene.
EUROC: EUROC dataset [9] contains different indoor and factory scenes in the real world. There are no dynamic objects but light intensity changes and violent shakings in it. It is used as a verification that our algorithm remains robust for static environments.
V-B Remove dynamic landmarks
After preprocessing to all landmarks during the feature tracking phase, the probabilistic model of landmarks’ minimum projection errors is closer to the chi-square distribution with one degree of freedom, and the coefficient of variation (CV) in every frame approaches to , as Fig.6 in VIODE dataset of day sequences from high to none. is set to be 4 and is set to be 1 in experiments.
After optimization based on sliding window using our visual residual for dynamic candidate landmarks and post-processing to them, confirmed dynamic landmarks will be removed from sliding window while others will be marked as static landmarks. The reduced dynamic visual residual result compared with original residual is displayed in Fig.5 and final labels of landmarks in Fig.3 is displayed in Fig.7.
V-C Qualitative Analyses
This subsection provides a quantitative analysis of our method, evaluating the localization accuracy of camera poses and the time taken. Tables I and II present the comparison of our proposed method with other state-of-the-art SLAM methods, using the RMSE (Root Mean Square Error) of ATE (Absolute Trajectory Error) and average tracking and optimization time cost as the evaluation metric. To comprehensively validate the fitness for environments within different dynamic ranges and the effectiveness of the proposed method, we conducted tests on the VIODE dynamic dataset and the EUROC static dataset.
Table I Comparison of Localization accuracy with state-of-the-art methods (RMSE of ATE in [m]). Noted that x represents failure case (diverged).
| Method | VIODE city | VIODE parking_lot | EUROC | ||||||||
| none | low | mid | high | none | low | mid | high | V1_03 | V2_03 | MH_03 | |
| VINS-Fusion | 0.158 | 0.169 | 0.193 | 0.211 | 0.103 | 0.126 | 0.149 | 0.166 | 0.216 | 0.352 | 0.290 |
| VINS-Fusion_ransac | 0.318 | 0.289 | 0.298 | 0.301 | 0.099 | 0.097 | x | 0.310 | 0.109 | 0.370 | 0.266 |
| Dyna-VINS | 0.157 | 0.140 | 0.149 | 0.122 | 0.115 | 0.099 | 0.126 | 0.124 | 0.175 | 0.479 | 0.349 |
| IDY-VINS | 0.148 | 0.145 | 0.135 | 0.126 | 0.092 | 0.089 | 0.111 | 0.088 | 0.143 | 0.229 | 0.276 |
The effects of the proposed method on the localization accuracy is analyzed as shown in Table I. And Table II shows the analysis of our method on the tracking and optimization time cost. The comparison of the trajectory results in the directions for different methods and ground truth are shown in Fig.8.
(1) The improvement of localization accuracy for different environments: According to Table I, our method has a better result than VINS-Fusion and Dyna-VINS while the RANSAC method has the problem of erroneously removing static inliers when facing large-scale dynamic scenes, like Fig.3(a), which leads to the decline and instability of the accuracy, although the RANSAC method can achieve good results in some cases. For outdoor sequences of VIODE from day_none to day_high, our method improves the localization accuracy by 6.3%, 14.2%, 30.1% and 40.3% compared to VINS-Fusion. And our method improves the localization accuracy by 10.7%, 29.4%, 25.5% and 47.0% compared to VINS-Fusion for indoor parking lot sequences of VIODE from parking_lot_none to parking_lot_high. The results of EUROC sequences in real world prove that our method still has a stable improvement effect on the localization accuracy in static environments. Although in some cases, the improvement effect is not as good as RANSAC method for EUROC dataset, but it is better than Dyna-VINS method. Moreover, the performance of Dyna-VINS method is related to complex parameters.
Table II Time cost experiment tested in day sequences of VIODE dataset from none to high. Noted that the max number of landmarks every frame is set to 120.
| Method |
| day_none | day_low | day_mid | day_high | day_none | day_low | day_mid | day_high | |||||
| VINS-Fusion | 19.562 | 20.671 | 20.473 | 19.568 | 39.565 | 38.894 | 39.245 | 37.642 | ||||
| VINS-Fusion_ransac | 19.621 | 21.704 | 21.752 | 22.155 | 38.393 | 38.984 | 39.848 | 38.887 | ||||
| Dyna-VINS | 19.353 | 19.504 | 19.498 | 19.703 | 55.755 | 48.155 | 43.628 | 49.859 | ||||
| IDY-VINS | 20.403 | 20.759 | 20.628 | 20.908 | 38.375 | 37.375 | 37.251 | 35.564 | ||||
(2) Better time cost performance: According to Table II, our method does not cause too much additional tracking time cost compared to RANSAC method. However, it should be noted that the time consumption will surely be higher than VINS-Fusion and Dyna-VINS which do not conduct additional elimination operations during the feature tracking phase. On the other hand, the optimized results converge more quickly for our method than VINS-Fusion and VINS-Fusion_ransac methods because of the elimination of obvious dynamic landmarks during feature tracking and the classification of dynamic candidate landmarks and static landmarks during optimization process. And the more dynamic objects there are in the environment, the better improvement on the consumption of optimization time will be for our method, while the RANSAC method often requires a longer optimization time to achieve convergence with a lower localization accuracy because of incorrect elimination. Meanwhile, Dyna-VINS method requires much more time for optimization due to its complex processing. The above two aspects of tracking time cost and optimization time cost demonstrate that our method generally has a better real-time performance.
VI CONCLUSIONS
In this paper, the robust visual-inertial motion prior SLAM system (IDY-VINS) has been proposed, which effectively handles dynamic landmarks using inertial motion prior for dynamic environments to varying degrees. It is a robust system that preprocesses potential dynamic landmarks and constructs bundle adjustment (BA) residual for dynamic candidate landmarks to reduce their impact. Experimental evidences demonstrate that our algorithm has a superior performance compared to other algorithms in both simulated and real environments, both dynamic and static environments.
In future work, we plan to extend the consideration of dynamic landmarks and candidate landmarks to the overall relocalization and loop closure detection processes of VI-SLAM.